[The following information applies to the questions displayed below.]

Causwell Company began 2021 with 14,000 units of inventory on hand. The cost of each unit was $4.00. During 2021 an additional 34,000 units were purchased at a single unit cost, and 24,000 units remained on hand at the end of 2021 (24,000 units therefore were sold during 2021). Causwell uses a periodic inventory system. Cost of goods sold for 2021, applying the average cost method, is $136,800. The company is interested in determining what cost of goods sold would have been if the FIFO or LIFO methods were used.

Required:
1. Determine the cost of goods sold for 2021 using the FIFO method. [Hint: Determine the cost per unit of 2021 purchases.] (Do not round intermediate calculations.)

2. Determine the cost of goods sold for 2021 using the LIFO method. (Do not round intermediate calculations.)
Answer #1

Average cost per unit = 136800/24000 = 5.7

5.7 = (14000*4+34000*X)/48000

273600 = 56000+34000X

X = 6.4

1) FIFO cost of goods sold = (14000*4+10000*6.4) = 120000

2) LIFO Cost of goods sold = 24000*6.4 = 153600